the aircraft has 86 more seats than the second aircraft. the third aircraft has 52 fewer seats than the second aircraft. if their total number of seats is 415, find the number of seats for each aircraft:
the first aircraft has __ seats. simplify your answer

Answer by rfer(16322)

x+86
x
x-52
x+86+x+x-52=415
3x+34=415
3x=415-34
x=381/3
x=127 2nd plane
x+86=213 1st plane
x-52=75 3rd plane